OpenStudy (anonymous):

Solve for x: 1 over 2 x + 2 = 22. 40 20 48 10

OpenStudy (anonymous):

is it \[\frac{1}{2x} +2\] or \[\frac{1}{2}x+2\] or \[\frac{1}{2x+2}\]